OpenStudy (alprincenofl):

True or False? The recursive definition of the set S={1,5,9,13,17,…} is 1∈S;x∈S→x+4∈S.

OpenStudy (alprincenofl):

Screenshot: http://prntscr.com/57m71r

OpenStudy (anonymous):

false

OpenStudy (anonymous):

False

OpenStudy (alprincenofl):

Why it's false? I think it's True. Both of "x" and "4" is a part of the set "S" so: x+4=?? Replace any number of the set with "x": 1+4=5 or 5+4=9 or etc all belong to the set "S" so it's True @nbandy0128 @Saquara

OpenStudy (alprincenofl):

@e.mccormick @jim_thompson5910 Can anyone please confirm the answer?

jimthompson5910 (jim_thompson5910):

That looks right, but I'm not familiar with this notation. So I'd get a 2nd opinion.

OpenStudy (alprincenofl):

@jim_thompson5910 What's your opinion please, true or false? I am confused

jimthompson5910 (jim_thompson5910):

it looks true, but again not 100% sure